# Verification of Geometric Robustness of Neural Networks via Piecewise Linear Approximation and Lipschitz Optimisation

### v1

@@ Abstract
We address the problem of verifying neural networks against geometric transformations of the input image, including rotation, scaling, shearing, and translation. The proposed method computes provably sound piecewise linear constraints for the pixel values by using sampling and linear approximations in combination with branch-and-bound Lipschitz optimisation. A feature of the method is that it obtains tighter over-approximations of the perturbation region than the present state-of-the-art. We report results from experiments on a comprehensive set of benchmarks. We show that our proposed implementation resolves more verification cases than present approaches while being more computationally efficient.
@@ 1 Introduction
Neural networks as used in mainstream applications - including computer vision - are known to be fragile and susceptible to adversarial attacks [19]. The area of formal verification of neural networks is concerned with the development of methods to establish whether a neural network is robust, with respect to its classification output, to variations of the image. A large body of literature has so far focused on norm-bounded input perturbations, aiming to demonstrate that imperceptible adversarial alterations of the pixels cannot alter the classifier’s classification ($\ell_{p}$ robustness). In safety-critical applications such as autonomous driving, however, resistance to norm-bounded perturbations is inadequate to guarantee safe deployment. In fact, image classifiers need to be robust against a number of variations of the image, including contrast, luminosity, hue, and beyond. A particularly important class of specifications concerns robustness to geometric perturbations of the input image [24, 28, 33, 2]. These may include translation, shearing, scaling, and rotation.
@@
The rest of this paper is organized as follows: Section 2 discusses related work. In Section 3 we introduce the problem of verifying neural networks against geometric robustness properties. In Section 4 we present our novel piecewise linear approximation strategy via sampling, optimisation and shifting. In Section 5 we discuss the experimental results obtained and contrast the present method against the state-of-the-art on benchmark networks. We conclude in Section 6. Our code is available at [1].
@@ 2 Related Work
Formal verification of geometric robustness
One of the earliest works [29] discretises the transformation domains, enabling robustness verification through the evaluation of the model at a finite number of discretised transformations. In contrast to Pei et al. [29], we here focus on continuous domains, which do not allow exhaustive evaluation. Previous work on continuous domains relies on over-approximations, whereby, for each pixel, the set of allowed values under the perturbation is replaced by a convex relaxation [24, 33, 2]. In particular, [24] and [33] use an $l_{\infty}$ norm ball and intervals respectively, resulting in loose over-approximations. Balunović et al. [2] devise more precise convex relaxations by computing linear approximations with respect to the transformation parameters. In this work, we further improve precision compared to Balunović et al. [2] by deriving piecewise linear approximations. While the above works consider the geometric transformation as a whole (see Section 3), Mohapatra et al. [28] decompose the transformation into network layers to be pre-pended to the network under analysis, resulting in looser approximations when using standard neural network verifiers. More recently, randomised smoothing techniques have been investigated for geometric robustness [15, 26, 16]: differently from our work, these only provide probabilistic certificates. Finally, Yang et al. [44] recently presented a method to train networks more amenable to geometric robustness verification. Our work is agnostic to the training scheme: we here focus on the more challenging general case.
@@ 3 Geometric robustness verification
Neural networks for classification
@@
Attack model via geometric transformation
@@ 4.3 Lipschitz optimisation for obtaining sound piecewise linear bounds
However, computing the exact maximum $\xi^{*}$ is computationally hard due to the nonconvexity, nonlinearity and nonsmoothness of $\underline{f}_{u,v}(\kappa)$. Instead, given any $\epsilon>0$, we can use a branch-and-bound Lipschitz optimisation procedure (see Algorithm 1) to find $\underline{\xi}^{*}\in\mathbb{R}$ satisfying $\underline{\xi}^{*}\leq\xi^{*}_{u,v}\leq\underline{\xi}^{*}+\epsilon.$
@@
This proof is motivated by [22]. In order to prove Proposition 3, we first state a useful result from [22, Lemma 3]. Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be Lipschitz continuous over an open set $\Omega\subset\mathbb{R}^{n}$. We denote $\mathrm{Diff}(\Omega)$ as the subset of $\Omega$ where $f(x)$ is differentiable. We also let $\mathcal{D}$ be the set of $(x,v)\in\mathbb{R}^{2n}$ for which the directional derivative, $\nabla_{v}f(x)$, exists and $x\in\Omega$. Finally, we let $\mathcal{D}_{v}$ be the set $\mathcal{D}_{v}=\{x\in\mathbb{R}^{n}\mid(x,v)\in\mathcal{D}\}.$ Then, we have the following inequality [22, Lemma 3]
@@
Furthermore, considering the inequality in (15[22, Lemma 3], we have
@@
Algorithm 1 Branch-and-bound Lipschitz Optimisation Procedure
Input: $\underline{f}_{u,v}$, $\mathcal{B}$, $\epsilon$, $n$, $N$
Output: $\xi^{*}$
5:     for $i\leftarrow 1$ to $N$ do
7:           $\mathcal{B}_{i,i+N}=\text{split}(\mathcal{B}_{i})$.
11:        end if
12:     end for
13:  end while
14:  return $\xi^{*}=\max_{i=1,\ldots,N}\underline{f}_{{\max},i}$.
Branch-and-bound Lipschitz optimisation procedure. Similar to [2], we use a branch-and-bound procedure - outlined in Algorithm 1 - where $\underline{f}_{u,v}$ and $\mathcal{B}$ are given as inputs alongside the Lipschitz error, $\epsilon$, and samples per subdomain, $n$. The procedure first samples the violation function $\underline{f}_{u,v}$, obtaining maximum value candidates, this is placed in a list of 3-tuples with the upper bound, $\underline{f}_{{\rm bound},i}$, and corresponding domain, $\mathcal{B}_{i}$. The key upper bound operation $\rm bound(\cdot)$ in Algorithm 1 is obtained using (14). We then check whether each 3-tuple in our list meets the termination criteria, as parameterised by $\epsilon$. If the requirement is satisfied for all elements then we terminate and return $\underline{\xi}^{*}$. Until the requirement is met for every list element we iteratively split unsatisfied subdomains. This process is repeated until a satisfactory maximum candidate is found, splitting $\kappa$ in each iteration. We can ignore any sub-domain, $\kappa_{\mathfrak{1}}^{n}$, of $\kappa_{\mathfrak{1}}$ where the function bound $\underline{f}_{\text{bound}}$ in $\kappa_{\mathfrak{1}}^{n}$ is smaller than a maximum value candidate $\underline{f}_{\text{max}}$ in any other sub-domain. Deciding how to split subdomains is non-trivial for higher dimensional parameter spaces. In the case $\kappa\in\mathbb{R}^{1}$ we need only decide where to split on a single axis; for which we use the domain midpoint. The crux of our algorithm is approximating the gradient of $\underline{f}_{u,v}(\kappa)$ when it is differentiable, as stated in Proposition 3 (see appendix for further details on the branch-and-bound procedure). For bounding the violation of piecewise linear bounds we can consider the piecewise bound itself to be made of $q$ linear sub-regions with each one bounded by the intersection with the neighbouring linear piece - or the lower and upper bounds on the transformation parameters. We can then bound the Lipschitz constant in the same way as for a single linear bound, instead starting with $q$ sub-domains. Solving the Lipschitz bounding procedure for each linear segment over only its local domain in this way enables us to bound the Lipschitz constant of a piecewise linear bound in the same time as we a linear bound takes.
@@ 5 Experimental Evaluation
In this section we present three sets of results: (i) a quantitative study directly comparing the model-agnostic bounds produced by our piecewise linear approach against the state-of-the-art linear bounds [2], (ii) an empirical evaluation of verification results obtained using linear and piecewise linear bounds, without input splitting and using the same neural network verifier [4], (iii) a comparison with state-of-the-art results previously reported in the literature [2].
@@ 5.1 Experimental setup
We consider the MNIST image recognition dataset [25]. In line with the previous literature [2], we use two fully-connected ReLU networks, MLP2 and MLP6, and one convolutional ReLU network, CONV, from the first competition for neural network verification (VNN-COMP) [38]. The fully-connected networks comprise 2 and 6 layers respectively. Each layer of each of the networks has 256 ReLU nodes. The convolutional network comprises two layers. The first layer has 32 filters of size $5\times 5$, a padding of 2 and strides of 2. The second layer has 64 filters of size of $4\times 4$, a padding of 2 and strides of 1. Additionally, we employ a larger convolutional ReLU network from relevant previous work [2], composed of three layers: a convolutional layer with 32 filters of size $4\times 4$ and strides of 2, a convolutional layer with 64 filters of size $4\times 4$ and strides of 2, and a fully connected layer with 200 nodes. All experiments were carried out on an Intel Core i9-10940X (3.30GHz, 28 cores) equipped with 256GB RAM and running Linux kernel 5.12.
@@ 5.2 Experimental results
PWL vs L: comparing areas
Figure 2 is a direct comparison of bound tightness between our piecewise linear bounds and the current state-of-the-art linear bounds [2]. For each image, linear and piecewise linear bounds are generated, each one capturing the reachable pixel values for a given transformation. We always use two piecewise segments ($q=2$) and use a Lipschitz error of 0.01 to compute bounds.The area enclosed by each set of bounds is then calculated and averaged for every pixel over all images. In each case the piecewise linear bounds are guaranteed to be tighter (enclose a smaller area) than the linear bounds, as in Section 4. Figure 2 shows the relative area (specifically, $1-(V_{\text{PWL}}/V_{\text{L}})$ with $V_{\text{PWL}}$ and $V_{\text{L}}$ being the volume enclosed by the piecewise linear and linear bounds, respectively) of the two bound types. In Figure 2, there is an initial increase in relative tightness for all transformations – this is a result of linear bounds being unable to efficiently capture the increasing nonlinearity in the pixel value curve, $\mathcal{G}_{u,v}(\kappa)$. After an initial increase, behaviour for different transformations diverges. For rotation, the relative advantage of the piecewise bounds continues to increase up to 15 degrees. For scaling, however, there is a peak at 1.25$\times$ magnification, followed by a decrease in the relative tightness. This result is explained by a corresponding increase in the complexity of the pixel value curve. Notably, the piecewise bounds are best suited to nonmonotonic pixel value curves with a single, sharp vertex. For curves with many vertices and large fluctuations, piecewise linear bounds become increasingly linear (the gradient of the pieces converge) to maintain convexity. Though this is the case for $q=2$, as we study here, for larger numbers of piecewise segments the advantage over linear bounds will continue to hold, as the piecewise bounds approximate the convex hull of the pixel values for $q\rightarrow\infty$The plots for shearing and translation show a similar pattern to scaling. Although the relative tightness may decrease for larger transformations, the total bounded area increases, making any proportional reduction in area more significant.
PWL vs L: verification results
Table 1 reports the experimental results obtained for verification queries using VENUS, on the VNN-COMP networks. For each type of input bound – piecewise linear and linear – the table shows the percentage of certified images (Verified column), the percentage of images for which a valid counter example was found (Falsified column), and the average verification time. We verify the robustness of each of the networks with respect to one of four transformations - rotation, scaling, shearing, or translation - on 50 randomly selected images from the MNIST test set. For each verification query we use a timeout of 30 minutes. We observe a considerable performance advantage using piecewise linear bounds for the convolution network, in every case, at least doubling the count of verifiable images. For the 6-layer MLP network, many of the transformations tried were unable to be verified, leading to numerous counter examples and time-outs. However, for every transformation the piecewise linear bounds were able to find more counter examples than linear bounds – this is a result of the improved tightness of piecewise linear bounds. For the 2-layer MLP, results across the bound types are very similar, in some cases they are equal. This is due to two factors, both of which stem from the network’s small size. Firstly, the 2-layer network is the least robust of all three. Accordingly, our results are for very small transformations for which the pixel value curve is approximately linear. In these cases, linear bounds can capture the input set as well as piecewise linear bounds. Secondly, the advantage of piecewise linear bounds’ tightness is compounded over each layer of a network – the 2-layer MLP is so small that this effect is minimal, further aligning the performance of the approaches. Finally, the use of piecewise linear constraints result in a reduction of average verification times on both the 6-layer MLP and the convolutional network: this is due to the fact that their relative tightness compensates for the additional cost of their encoding, leading the employed MILP-based verifier to positive lower bounds on the verification problem (1) in less time.
Comparison with literature results
@@
In Table 2 we provide a comparison of verification results obtained using VENUS with both linear and piecewise linear constraints, with the DeepG [2] results, obtained using linear constraints and the DeepPoly [33] verifier, which relies on a relatively loose LP relaxation of (1). Further, we use a MILP-based verifier which enabled us to add the pixel domain constraints in addition to our transformation-based bounds. This, coupled with the tighter verifier, enables our linear bounds to out-perform those from DeepG. We consider a MNIST benchmark presented in Balunović et al. [2], where a 30 degree rotation transformation is verified by way of 10, 3-degree sub-problems. This is in contrast to Table 1, where each perturbation is represented by a single set of bounds and a single verifier call per image. Table 2 shows that, even under the small-perturbation setting, the use of tighter verification algorithms (L versus DeepG) increases the number of verified properties. Furthermore, we show that the method proposed in this work, PWL, leads to the tightest certification results. Nevertheless, we point out that verifying perturbations through a series of sub-problems is extremely expensive, as it requires repeated calls to both neural network verifiers, and to the constraint-generation procedure (including the branch-and-bound-based Lipschitz optimisation). For this reason, we focus on verification the setting without transformation splitting, and aim to maximize certifications through the use of tight verifiers and over-approximations of the geometric transforms.

### v2

@@ Abstract
We address the problem of verifying neural networks against geometric transformations of the input image, including rotation, scaling, shearing, and translation. The proposed method computes provably sound piecewise linear constraints for the pixel values by using sampling and linear approximations in combination with branch-and-bound Lipschitz optimisation. The method obtains provably tighter over-approximations of the perturbation region than the present state-of-the-art. We report results from experiments on a comprehensive set of verification benchmarks on MNIST and CIFAR10. We show that our proposed implementation resolves up to 32% more verification cases than present approaches.
@@ 1 Introduction
Neural networks as used in mainstream applications, including computer vision, are known to be fragile and susceptible to adversarial attacks [18]. The area of formal verification of neural networks is concerned with the development of methods to establish whether a neural network is robust, with respect to its classification output, to variations of the image. A large body of literature has so far focused on norm-bounded input perturbations, aiming to demonstrate that imperceptible adversarial alterations of the pixels cannot alter the classifier’s classification ($\ell_{p}$ robustness). In safety-critical applications such as autonomous driving, however, resistance to norm-bounded perturbations is inadequate to guarantee safe deployment. In fact, image classifiers need to be robust against a number of variations of the image, including contrast, luminosity, hue, and beyond. A particularly important class of specifications concerns robustness to geometric perturbations of the input image [23, 28, 33, 1]. These may include translation, shearing, scaling, and rotation.
@@
The rest of this paper is organized as follows: Section 2 discusses related work. In Section 3 we introduce the problem of verifying neural networks against geometric robustness properties. In Section 4 we present our novel piecewise linear approximation strategy via sampling, optimisation and shifting. In Section 5 we discuss the experimental results obtained and contrast the present method against the state-of-the-art on benchmark networks. We conclude in Section 6. Our code is publicly available on GitHub [^1] https://github.com/benbatten/PWL-Geometric-Verification and a full version of the paper with its appendix is available on arXiv [^2] https://arxiv.org/abs/2408.13140.
@@ 2 Related Work
Formal verification of geometric robustness.
One of the earliest works [29] on this subject discretises the transformation domains, enabling robustness verification through the evaluation of the model at a finite number of discretised transformations. In contrast to Pei et al. [29], we here focus on continuous domains, which do not allow exhaustive evaluation. Previous work on continuous domains relies on over-approximations, whereby, for each pixel, the set of allowed values under the perturbation is replaced by a convex relaxation [23, 33, 1]. In particular, [23] and [33] use an $l_{\infty}$ norm ball and intervals respectively, resulting in loose over-approximations. Balunović et al. [1] devise more precise convex relaxations by computing linear approximations with respect to the transformation parameters. In this work we further improve precision over Balunović et al. [1] by deriving piecewise linear approximations. While the above works consider the geometric transformation as a whole (see Section 3), Mohapatra et al. [28] decompose the transformation into network layers to be pre-pended to the network under analysis, resulting in looser approximations when using standard neural network verifiers. More recently, randomised smoothing techniques have been investigated for geometric robustness [14, 26, 15]: differently from our work, these only provide probabilistic certificates. Finally, Yang et al. [44] recently presented a method to train networks more amenable to geometric robustness verification. Our work is agnostic to the training scheme: we here focus on the more challenging general case.
@@ 3 Geometric robustness verification
Neural networks for classification.
@@
Attack model via geometric transformation.
@@ 4.3 Lipschitz optimisation for obtaining sound piecewise linear bounds
However, computing the exact maximum $\xi^{*}$ is computationally hard due to the nonconvexity, nonlinearity and nonsmoothness of $\underline{f}_{u,v}(\kappa)$. Instead, given any $\epsilon>0$, we can use a branch-and-bound Lipschitz optimisation procedure to find $\underline{\xi}^{*}\in\mathbb{R}$ satisfying $\underline{\xi}^{*}\leq\xi^{*}_{u,v}\leq\underline{\xi}^{*}+\epsilon.$
@@
This proof is motivated by [21]. In order to prove Proposition 3, we first state a useful result from [21, Lemma 3]. Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be Lipschitz continuous over an open set $\Omega\subset\mathbb{R}^{n}$. We denote $\mathrm{Diff}(\Omega)$ as the subset of $\Omega$ where $f(x)$ is differentiable. We also let $\mathcal{D}$ be the set of $(x,v)\in\mathbb{R}^{2n}$ for which the directional derivative, $\nabla_{v}f(x)$, exists and $x\in\Omega$. Finally, we let $\mathcal{D}_{v}$ be the set $\mathcal{D}_{v}=\{x\in\mathbb{R}^{n}\mid(x,v)\in\mathcal{D}\}.$ Then, we have the following inequality [21, Lemma 3]
@@
Furthermore, considering the inequality in (15[21, Lemma 3], we have
@@
Branch-and-bound Lipschitz optimisation procedure. Similar to [1], we use a branch-and-bound procedure (See Appendix) where $\underline{f}_{u,v}$ and $\mathcal{B}$ are given as inputs alongside the Lipschitz error, $\epsilon$, and samples per subdomain, $n$. The procedure first samples the violation function $\underline{f}_{u,v}$, obtaining maximum value candidates, this is placed in a list of 3-tuples with the upper bound, $\underline{f}_{{\rm bound},i}$, and corresponding domain, $\mathcal{B}_{i}$. The key upper bound operation $\rm bound(\cdot)$ is obtained using (14). We then check whether each 3-tuple in our list meets the termination criteria, as parameterised by $\epsilon$. If the requirement is satisfied for all elements then we terminate and return $\underline{\xi}^{*}$. Until the requirement is met for every list element we iteratively split unsatisfied subdomains. This process is repeated until a satisfactory maximum candidate is found, splitting $\kappa$ in each iteration. We can ignore any sub-domain, $\kappa_{\mathfrak{1}}^{n}$, of $\kappa_{\mathfrak{1}}$ where the function bound $\underline{f}_{\text{bound}}$ in $\kappa_{\mathfrak{1}}^{n}$ is smaller than a maximum value candidate $\underline{f}_{\text{max}}$ in any other sub-domain. Deciding how to split subdomains is non-trivial for higher dimensional parameter spaces. In the case $\kappa\in\mathbb{R}^{1}$ we need only decide where to split on a single axis; for which we use the domain midpoint. The crux of our algorithm is approximating the gradient of $\underline{f}_{u,v}(\kappa)$ when it is differentiable, as stated in Proposition 3 (see appendix for further details on the branch-and-bound procedure). For bounding the violation of piecewise linear bounds we can consider the piecewise bound itself to be made of $q$ linear sub-regions with each one bounded by the intersection with the neighbouring linear piece - or the lower and upper bounds on the transformation parameters. We can then bound the Lipschitz constant in the same way as for a single linear bound, instead starting with $q$ sub-domains. Solving the Lipschitz bounding procedure for each linear segment over only its local domain in this way enables us to bound the Lipschitz constant of a piecewise linear bound in the same time as a linear bound takes.
@@ 5 Experimental Evaluation
In this section we present three sets of results: (i) a quantitative study directly comparing the model-agnostic bounds produced by our piecewise linear approach against the state-of-the-art linear bounds [1], (ii) an empirical evaluation of verification results obtained using linear and piecewise linear bounds, without input splitting and using the same neural network verifier [3], and (iii) a comparison of our results against the present state-of-the-art method [1].
@@ 5.1 Experimental setup
We consider the MNIST image recognition dataset [25] and CIFAR10 [24]. In line with the previous literature [1], we use two fully-connected ReLU networks, MLP2 and MLP6, and one convolutional ReLU network, CONV, from the first competition for neural network verification (VNN-COMP) [38]. The fully-connected networks comprise 2 and 6 layers respectively. Each layer of each of the networks has 256 ReLU nodes. The convolutional network comprises two layers. The first layer has 32 filters of size $5\times 5$, a padding of 2 and strides of 2. The second layer has 64 filters of size of $4\times 4$, a padding of 2 and strides of 1. Additionally, we employ a larger convolutional ReLU network from relevant previous work [1], composed of three layers: a convolutional layer with 32 filters of size $4\times 4$ and strides of 2, a convolutional layer with 64 filters of size $4\times 4$ and strides of 2, and a fully connected layer with 200 nodes. All experiments were carried out on an Intel Core i9-10940X (3.30GHz, 28 cores) equipped with 256GB RAM and running Linux kernel 5.12. DeepG experiment ANS: we do not use GPU in these experiments.
@@ 5.2 Experimental results
PWL vs L: comparing areas.
Figure 2 is a direct comparison of bound tightness between our piecewise linear bounds and the current state-of-the-art linear bounds [1]. For each image, linear and piecewise linear bounds are generated, each one capturing the reachable pixel values for a given transformation. We always use two piecewise segments ($q=2$) and use a Lipschitz error of 0.01 to compute bounds. The area enclosed by each set of bounds is then calculated and averaged for every pixel over all images. In each case the piecewise linear bounds are guaranteed to be tighter (enclose a smaller area) than the linear bounds, as in Section 4. Figure 2 shows the relative area (specifically, $1-(V_{\text{PWL}}/V_{\text{L}})$ with $V_{\text{PWL}}$ and $V_{\text{L}}$ being the volume enclosed by the piecewise linear and linear bounds, respectively) of the two bound types. In Figure 2, there is an initial increase in relative tightness for all transformations – this is a result of linear bounds being unable to efficiently capture the increasing nonlinearity in the pixel value curve, $\mathcal{G}_{u,v}(\kappa)$. After an initial increase, the behaviour for different transformations diverges. For rotation, the relative advantage of the piecewise bounds continues to increase up to 15 degrees. For scaling, however, there is a peak at 1.25$\times$ magnification, followed by a decrease in the relative tightness. This result is explained by a corresponding increase in the complexity of the pixel value curve. Notably, the piecewise bounds are best suited to nonmonotonic pixel value curves with a single, sharp vertex. For curves with many vertices and large fluctuations, piecewise linear bounds become increasingly linear (the gradient of the pieces converge) to maintain convexity. Though this is the case for $q=2$, as we study here, for larger numbers of piecewise segments the advantage over linear bounds will continue to hold, as the piecewise bounds approximate the convex hull of the pixel values for $q\rightarrow\infty$. The plots for shearing and translation show a similar pattern to scaling. Although the relative tightness may decrease for larger transformations, the total bounded area increases, making any proportional reduction in area more significant.
PWL vs L: verification results.
Table 1 reports the experimental results obtained for verification queries using VENUS, on the VNN-COMP networks. For each type of input bound – piecewise linear and linear – the table shows the percentage of certified images (Verified column), the percentage of images for which a valid counter example was found (Falsified column), and the average verification time. We verify the robustness of each of the networks with respect to one of four transformations - rotation, scaling, shearing, or translation - on 50 randomly selected images from the MNIST test set. For each verification query we use a timeout of 30 minutes. We observe a considerable performance advantage using piecewise linear bounds for the convolution network, in every case, at least doubling the count of verifiable images. For the 6-layer MLP network, many of the transformations tried could not be verified, leading to numerous counter examples and time-outs. However, for every transformation the piecewise linear bounds were able to find more counter examples than linear bounds – this is a result of the improved tightness of piecewise linear bounds. For the 2-layer MLP, results across the bound types are very similar, in some cases they are equal. This is due to two factors, both of which stem from the network’s small size. Firstly, the 2-layer network is the least robust of all three. Accordingly, our results are for very small transformations for which the pixel value curve is approximately linear. In these cases, linear bounds can capture the input set as well as piecewise linear bounds. Secondly, the advantage of piecewise linear bounds’ tightness is compounded over each layer of a network – the 2-layer MLP is so small that this effect is minimal, further aligning the performance of the approaches. Finally, the use of piecewise linear constraints result in a reduction of average verification times on both the 6-layer MLP and the convolutional network: this is due to the fact that their relative tightness compensates for the additional cost of their encoding, leading the employed MILP-based verifier to positive lower bounds on the verification problem (1) in less time.
Comparison with literature results.
@@
In Table 2 we provide a comparison of verification results obtained using VENUS with both linear and piecewise linear constraints, with the DeepG [1] results, obtained using linear constraints and the DeepPoly [33] verifier, which relies on a relatively loose LP relaxation of (1). Further, we use a MILP-based verifier which enabled us to add the pixel domain constraints in addition to our transformation-based bounds. This, coupled with the tighter verifier, enables our linear bounds to out-perform those from DeepG. We consider MNIST and CIFAR10 benchmark presented in Balunović et al. [1]. The MNIST example consists of verifying a 30 degree rotation transformation by way of 10, 3-degree sub-problems. This is in contrast to Table 1, where each perturbation is represented by a single set of bounds and a single verifier call per image. Table 2 shows that, even under the small-perturbation setting, the use of tighter verification algorithms (L versus DeepG) increases the number of verified properties. Furthermore, we show that the method proposed in this work, PWL, leads to the tightest certification results. The CIFAR10 example comprises a composition of rotation and shearing of 2 degrees and 2% respectively. This query is solved via 4 sub-problems (with each transformation domain split in half). The results show a 12% improvement for the PWL bounds over the DeepG result. However, much of this gain comes from the verifier itself. The gap between the linear bounds and their piecewise counterpart is 1%. We attribute this smaller gap to the relatively small domain over which each sub-problem runs. Nevertheless, we point out that verifying perturbations through a series of sub-problems is extremely expensive, as it requires repeated calls to both neural network verifiers, and to the constraint-generation procedure (including the branch-and-bound-based Lipschitz optimisation). For this reason, we focus on verification the setting without transformation splitting, and aim to maximize certifications through the use of tight verifiers and over-approximations of the geometric transforms.